OpenStudy (anonymous):

Is anyone willing to help with some 8th grade math?

OpenStudy (solomonzelman):

Sweet, we got some invisible mathematics on here:) lol, what is your question ?

OpenStudy (anonymous):

Consider the following pair of equations: y = –2x + 8 y = x – 1 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form.

OpenStudy (anonymous):

I have 7 questions

OpenStudy (solomonzelman):

Lets do one question per post. Is that okay?

OpenStudy (anonymous):

okay

OpenStudy (solomonzelman):

so we will go straight to your problem (that you posted). In front of you is the system of equations, and the solution is the intersection point. Which methods have you learned/like substitution, elimination, (graphing - if that is permitted) and/or matrix?

OpenStudy (solomonzelman):

I would do a direct substitution in this case.

OpenStudy (anonymous):

I dont know how to do that

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y=-2x+8 }\) \(\large\color{black}{ \displaystyle y=x-1 }\) so you can use your second equation. it tells you that y is same thing as x-1 so, instead of saying \(\large\color{black}{ \displaystyle \color{red}{y}=-2x+8 }\) , you can write: \(\large\color{black}{ \displaystyle \color{red}{x-1}=-2x+8 }\)

OpenStudy (solomonzelman):

Fair enough ?

OpenStudy (anonymous):

okay so I would type x - 1 + -2x +8 ?

OpenStudy (solomonzelman):

yes (don't leave your signs out though) \(\large\color{black}{ \displaystyle \color{red}{x-1}=-2x+8 }\)

OpenStudy (solomonzelman):

can you solve for \(\large\color{black}{ \displaystyle x}\) in this equation ?

OpenStudy (anonymous):

I dont know how to? I was gone during this lesson..

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle x-1=-2x+8 }\) \(\small\color{black}{ \displaystyle \bullet }\) Add \(\normalsize\color{black}{ \displaystyle 2x }\) to both sides. \(\small\color{black}{ \displaystyle \bullet }\) Add \(\normalsize\color{black}{ \displaystyle 1 }\) to both sides. \(\small\color{black}{ \displaystyle \bullet }\) Divide by \(\normalsize\color{black}{ \displaystyle 3 }\) on both sides.

OpenStudy (solomonzelman):

lets start with adding 2x to both sides. YOU HAD: \(\large\color{black}{ \displaystyle x-1=-2x+8 }\) NOW ADDING: \(\large\color{black}{ \displaystyle x-1\color{blue}{+2x}=-2x+8\color{blue}{+2x} }\) what will you get after doing this, can you tell me?

OpenStudy (anonymous):

x-3x=8x?

OpenStudy (solomonzelman):

I apologize, but you aren't showing enough background for solving this problem. I would strongly advise to get assistance with (and so respectively): 1) What "like terms" are and how to add them. 2) Substitution and Elimination methods of solving a system of equations. to proceed without knowing at least the #1, would be irrational.